Complete the following​ table:  ​(Enter your responses rounded to two decimal​ places.)

Initial value

New value

Percentage change

7272

116116

nothing​%

160160

122122

nothing​%

192192

nothing

Help!

LOADING...

9.389.38​%

5 comma 6305,630

11 comma 64111,641

nothing​%
Answer #1

(1) Initial value = 72, New value = 116

% Change = (116/72) - 1 = 1.6111 - 1 = 0.6111 = 61.11%

(2) Initial value = 160, New value = 122

% Change = (122/160) - 1 = 0.7625 - 1 = -0.2375 = -23.75%

(3) Initial value = 192, % Change = 9.38% = 0.0938

If New value be V, then

(V/192) - 1 = 0.0938

V/192 = 1.0938

V = 210.01

(4) Initial value = 5,630 and New value = 11,641

% Change = (11,641/5,630) - 1 = 2.0677 - 1 = 1.0677 = 106.77%